OpenStudy (anonymous):

Which of the following equations has no solution? 3x + 1 = 2x + 1 3x + 1 = 3x + 1 3x + 1 = 3x + 2

OpenStudy (anonymous):

b

OpenStudy (anonymous):

sorry c

OpenStudy (anonymous):

Equation (1): 3x + 1 = 2x + 1 x + 1 = 1 x = 0 This is valid. Equation (2): 3x + 1 = 3x + 1 0 + 1 = 1 0 = 0 This is valid ('all solutions' we call this) Equation (3): 3x + 1 = 3x + 2 0 + 1 = 2 0 = 1 is not a valid statement, so no solution!

OpenStudy (kla01):

C has no solution.

OpenStudy (anonymous):

the answer was C